Given the function g(x) = 8x − 2, compare and contrast g(−2) and g(4). Choose the statement that is true concerning these two values.

Answer by richard1234(7193)

g(-2) = 8(-2) - 2 = -18, g(4) = 8(4) - 2 = 30

What statements are there to choose from?